gcc8617:

what is 10 to the second power

Angle:

\(10^2 = 10 \times 10 = ~?\)

gcc8617:

yea

Angle:

what is 10 times 10 ?

narutouzamaki5:

huh well then whats  its multiply

narutouzamaki5:

its a method called lattice very helpful

narutouzamaki5:

u go across and multiply so whats 0 times 1

Renne:

it's like if u were to go by ten's ten times what would u stop at?

DumbFounded:

you could just multiply 10 times 10

DumbFounded:

so 82 = 8 × 8 = 64. In words: 82 could be called "8 to the power 2" or "8 to the second power", or simply "8 squared".

Renne:

Yeah and what would your answer be?

Renne:

dont give out the answer tho u have to let them solve it.

Renne:

your going to get a warning for giving the answer

Eli9lizzy:

howcome u cant say the answers

DumbFounded:

how ever many little numbers you have next to the big number you multiply the big number by itself that many time

Renne:

they want the person who posted the qestion to fugier it out

Eli9lizzy:

oh ok

DumbFounded:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @DumbFounded how ever many little numbers you have next to the big number you multiply the big number by itself that many time \(\color{#0cbb34}{\text{End of Quote}}\) @gcc8617

Angle:

guys, enough is enough - let the original person respond and we're done here it is unnecessary to say 10 times 10 so many different ways

Perla112236:

just multiply 10x10

narutouzamaki5:

im ghelping him by dm ok

Renne:

Ok.

gcc8617:

u mean her

Renne:

Hopefully it lets u

DumbFounded:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @gcc8617 u mean her \(\color{#0cbb34}{\text{End of Quote}}\) do you understand the information being handed too you?

Renne:

I think she does.

Renne:

I hope she does

Eli9lizzy:

i dmed her too

Renne:

ok